# Online and Offline Space-Filling Input Design for Nonlinear System Identification: A Receding Horizon Control-Based Approach

### v1

@@ Abstract
The space-filling performance of the proposed strategy is evaluated on an artificial example and compared to state-of-the-art methods, demonstrating superior efficiency in exploring process operating spaces.
@@ I INTRODUCTION
The primary objective of input design methodologies is to acquire precise and comprehensive information about the process behavior intended to be modeled. This is particularly challenging when capturing nonlinearities, which require data spanning the entire operational range [4], in combination with dynamic processes, whose operating points cannot be manipulated instantaneously and thus require both static and transient behaviors to be represented in the dataset [5]. These difficulties are further exacerbated by real-world constraints such as limited time, process constraints, and high measurement costs [6].
@@ I-A Localization in Related Work
Model-based methods can be further subdivided into two primary categories: (i) Approaches that aim to minimize the variance of the parameter estimate by leveraging the Fisher information matrix (FIM) [7, 8, 9]. However, these methods are fundamentally limited by the assumption of an unbiased estimator, as required by the Cramér-Rao bound [10]. In practice, this assumption is often violated due to limited prior knowledge about the process. (ii) Approaches, typically following an active learning strategy, which aim to minimize the (epistemic) uncertainty of the underlying model [11, 12, 13]. Model-based input signal design can, in principle, generate highly informative datasets by directly optimizing process-focused optimality criteria. However, the success of these approaches heavily relies on the model‘s capability to accurately capture the process. Ensuring this capability in turn depends on prior knowledge – information that is often unavailable.
@@ I-B Space-Filling Design in Model Input Space
The proposed method further distinguishes from existing approaches by introducing a novel optimality criterion embedded within a receding horizon control (RHC)-based optimization framework. In essence, the approach aims to generate a space-filling design within the relevant operating region of a surrogate model’s input space. This input space may correspond, for instance, to the joint input-state space, as in [18, 19, 20], or to the model regressor space, as considered in [5, 16, 17]. The fundamental premise is that the information an excitation signal provides about a process is inherently tied to the distribution it induces within the selected model space. Thus, if all relevant regions of a model’s input space are sufficiently represented in the estimation dataset, the resulting model can capture accurately the desired process behavior. This consideration is particularly relevant for Markovian processes, where future states can be fully predicted given the knowledge of the current state [21]. Consequently, the task of acquiring valuable information about the process behavior can be reframed as the challenge of sufficiently exciting the regions of interest within the model input space.
The generalization error, i.e., model errors in regions of the operating space not covered by the estimation data, is one of the main sources of inaccuracies in nonlinear system identification [1, 2]. A space-filling design within the operating region of interest ensures comprehensive data collection of the process behavior intended to be modeled, minimizing knowledge gaps. Moreover, evenly distributed data mitigates data bias toward specific local operating regions, reducing the risk of overfitting. Consequently, space-filling excitation promotes the development of models that generalize well, contributing to robust and reliable predictions.
@@ II PROBLEM FORMULATION
$\hat{y}(k)=f\bigl{(}\underline{x}(k)\bigr{)}\hskip 3.0pt\quad\forall\,k\in\mathbb{N}\ ,$ (2)
@@
$\begin{split}{\underline{x}}^{T}(k)=&[u_{1}(k-1),\ldots,u_{1}(k-m),\ldots,\\ &u_{n_{u}}(k-1),\ldots,u_{n_{u}}(k-m),\\ &y_{1}(k-1),\ldots,y_{1}(k-m),\ldots,\\ &y_{n_{y}}(k-1),\ldots,y_{n_{y}}(k-m)]\\ &\forall k\in\mathbb{N}_{0}\ ,\\ \end{split}$ (3)
@@
In a data-driven modeling framework, (2) is inferred from estimation data, making its accuracy contingent on the quality of the available dataset. The general objective of input design is to generate a dataset
$\begin{split}{\underline{U}}^{T}=[\underline{u}(1),\ldots,\underline{u}(N)]\ ,\hskip 3.0ptN\in\mathbb{N}\ ,\end{split}$ (4)
of size $N$, where $\underline{u}^{T}(k)=[{u}_{1}(k),\ldots,{u}_{n_{u}}(k)]\ \forall k\in\{1,2,\ldots,N\}$, which should capture as much information as possible about the process behavior intended to be modeled. This ensures that the function in (2) can be estimated accurately, thereby minimizing the error in (1). In this contribution, the objective is pursued by optimizing the surrogate model’s input space distribution according to a space-filling optimality criterion. For an NARX structure, this distribution corresponds to the regressor space distribution, which can be expressed for an excitation dataset as
$\begin{split}\underline{{X}}^{T}=[{\underline{x}}(1),{\underline{x}}(2),\ldots,{\underline{x}}(N)]\ ,\hskip 3.0ptN\in\mathbb{N}\ \ .\end{split}$ (5)
The initial state of the process $\underline{x}(0)$, as well as the safe operating region for the controllable inputs $\mathcal{U}$ and the regressor space $\mathcal{X}$ are assumed to be known. Therefore, when generating a space-filling regressor space distribution, the following constraints must be satisfied:
$\begin{split}&\underline{{u}}(k)\in\mathcal{U}\hskip 3.0pt\quad\forall\,k=\{1,2,\ldots,N\}\ ,\\ &\underline{{x}}(k)\in\mathcal{X}\hskip 3.0pt\quad\forall\,k=\{0,1,\ldots,N\}\ .\end{split}$ (6)
@@ III RHC-BASED SPACE-FILLING DESIGN
In this section, the proposed input signal design methodology is introduced and its key contributions are highlighted. First, we develop a receding horizon control (RHC)-based iterative optimization procedure, following a similar approach to [22, 23]. Second, we formulate a novel optimality criterion that is embedded in the RHC-strategy and optimizes a surrogate model’s input space distribution to achieve a space-filling design within the operating region of interest, thus ensuring comprehensive and balanced information acquisition about the process behavior intended to be modeled.
@@ III-A Novel Optimality Criterion
$\begin{split}\underline{{\psi}}(j)\in\mathcal{C}\hskip 3.0pt\quad&\forall j={1,2,\ldots,N_{\Psi}}\\ &N_{\Psi}\in\mathbb{N}\ ,\end{split}$ (7)
@@
$\begin{split}J(\underline{\hat{X}},\underline{{\Psi}})&=\sum_{j=1}^{N_{\Psi}}d_{\mathrm{NN}}\bigl{(}{\underline{\psi}}(j),\underline{\hat{{X}}}\bigr{)}\\ \mathrm{with\,}\ \ &d_{\mathrm{NN}}\bigl{(}{\underline{\psi}}(j),\underline{\hat{{X}}}\bigr{)}=\underset{1\leq i\leq N}{\min}||\underline{\hat{x}}(i)-\underline{\psi}(j)||_{2}\ .\end{split}$ (8)
Since an NARX structure is chosen in this contribution, the input space distribution of the surrogate model corresponds to the regressor space distribution $\hat{\underline{X}}$, which is approximated by the surrogate model. This approximation is necessary since access to future outputs – required for executing the RHC-based optimization procedure, see Sec. III-B – as well as past outputs when running the proposed algorithm offline, is essential but unavailable. Therefore, the yet unknown true process outputs $y$ are substituted with the surrogate model’s predictions $\hat{y}$. For a first-order dynamic system ($m=1$) with a single input ($n_{u}=1$) and a single output ($n_{y}=1$), the surrogate model constructs the regression vector as follows:
$\underline{x}^{T}(k)=[u(k-1),\hat{y}(k-1)]\quad\forall k\in\mathbb{N}_{0}\ .$ (9)
Accordingly, the surrogate regressor space distribution used for computing the optimality criterion in (8) is given by
$\begin{split}\underline{{\hat{X}}}^{T}=[{\underline{\hat{x}}}(1),{\underline{\hat{x}}}(2),\ldots,{\underline{\hat{x}}}(N)]\ ,\hskip 3.0ptN\in\mathbb{N}\ \ .\end{split}$ (10)
A key advantage of the proposed optimality criterion is its piecewise $C^{1}$ differentiability, as well as its $C^{0}$ differentiability, which facilitates efficient optimization using gradient-based techniques.
@@ III-B RHC-Based Optimization Procedure
The core principle of the RHC-based procedure lies in its iterative optimization. In each iteration, the optimal sequence $\hat{\underline{U}}_{\mathrm{opt}}$ over a finite prediction horizon $L$ is determined as follows:
[…]
Here, $\mathcal{M}_{\theta}$ represents the surrogate model, $\underline{U}$ denotes the previously optimized input data points, and $\underline{\hat{U}}$ refers to the input data points yet to be optimized. To provide further clarification, the regressor space constructed using the surrogate model for the case $m=1,n_{u}=1,n_{y}=1$ is defined by
$\underline{\hat{X}}$ $=\begin{bmatrix}u(1)&\hat{y}(1)\\ \vdots&\vdots\\ u(k-1)&\hat{y}(k-1)\\ \hat{u}(k)&\hat{y}(k)\\ \vdots&\vdots\\ \hat{u}(k+L-1)&\hat{y}(k+L-1)\end{bmatrix}\ ,$ (12)
@@
Once $\underline{\hat{U}}_{\mathrm{opt}}$ is determined, a forward time shift is performed, updating $k\rightarrow k+1$. This optimization procedure is iteratively repeated until the entire signal is generated. Constraining the controllable inputs and the surrogate regressor space, as outlined in Eq. (LABEL:eq:constraints), is particularly beneficial for real-world applications, as it prevents operation in undesirable regions.
In this contribution, the optimization focuses on the surrogate regressor space distribution of the generated excitation signal. However, optimizing the joint input-state space obtained by a surrogate, as explored in [18, 19, 20], is also a viable approach. Pseudocode for the proposed procedure is presented in Algorithm 1.
xxxxxx
@@
Initialization: Safe operating region of the controllable inputs $\mathcal{U}$, safe operating region of the regressor space $\mathcal{X}$, supporting points $\underline{\Psi}$, surrogate model $\mathcal{M}_{\theta}$, initial model regressor $\underline{\hat{x}}(0)$, dimension $p$ of the regressor space distribution.
@@
     $\underline{\hat{U}}_{\mathrm{opt}}=\underset{\underline{\hat{U}}}{\mathrm{arg\,min}}\ J\bigl{(}\underline{\hat{X}},\underline{{\Psi}}\bigr{)}$triangleright Get optimal xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxsequence.
$\begin{split}\mathrm{with}\ &\underline{\hat{X}}=\mathcal{M}_{\theta}(\underline{\hat{U}},\underline{U},\underline{\hat{x}}(0),T_{s})\\ \mathrm{s.\,t.\ }\ &\underline{\hat{u}}(j)\in\mathcal{U}\hskip 3.0pt\quad\forall\,j=\{k,k+1,\ldots,k+L-1\}\\ &\underline{{u}}(j)\in\mathcal{U}\hskip 3.0pt\quad\forall\,j=\{1,2,\ldots,k-1\}\\ &\underline{\hat{x}}(j)\in\mathcal{\mathcal{X}}\hskip 3.0pt\quad\forall\,j=\{0,1,\ldots,k,\ldots,k+L-1\}\\ &\underline{\psi}(j)\in\mathcal{C}\quad\forall\,j=\{1,2,\ldots,N_{\Psi}\}\\ \end{split}$
     Apply $\underline{U}\leftarrow\underline{{U}}\cup\underline{\hat{u}}_{\mathrm{opt}}(k)$. $\triangleright$ Append optimal xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx data point.
@@ IV EVALUATION
Unfortunately, the concept of space-filling design is not well-defined, particularly for dynamic processes. However, two key aspects are generally considered important: (i) ensuring that no region remains underrepresented, meaning that no gaps are left [24], and (ii) achieving a distribution that covers the considered area as uniformly as possible [25]. Hence, to evaluate the space-filling properties, two metrics are applied. First, we examine the radius of the largest ball within the region of interest $\mathcal{C}$, defined as
@@
where $X$ denotes the regressor space probability distribution, $E$ the uniform distribution, ${M}=\frac{1}{2}({X}+{E})$ the mixture distribution, and $D_{\mathrm{KL}}$ the Kullback-Leibler divergence [26].
Furthermore, both execution approaches presented–namely, online execution with continuous updates of the model parameters based on current process observations, and offline execution with a fixed model, utilizing only an approximate process time constant as in [5, 16, 17]–are considered. To this end, a Hammerstein process is employed as example process in the evaluation, defined by:

### v2

@@ Abstract
The space-filling performance of the proposed strategy is evaluated on an artificial example and compared to state-of-the-art methods, demonstrating superior efficiency in exploring process operating regions.
@@ I Introduction
The primary objective of input design methodologies is to acquire precise and comprehensive information about the process behavior intended to be modeled. This is particularly challenging when capturing nonlinearities, which require data spanning the entire operational range [4], in combination with dynamic processes, whose operating points cannot be manipulated instantaneously and thus require both static and transient behaviors to be represented in the dataset [5]. These difficulties are further exacerbated by real-world constraints, such as limited time, process constraints, and high measurement costs [6].
@@ I-A Localization in Related Work
Model-based methods can be further subdivided into two primary categories: (i) Approaches that aim to minimize the variance of the parameter estimate by leveraging the Fisher information matrix (FIM) [7, 8, 9]. However, these methods are fundamentally limited by the assumption of an unbiased estimator, as required by the Cramér-Rao bound [10]. In practice, this assumption is often violated due to limited prior knowledge about the process. (ii) Approaches, typically following an active learning strategy, which aim to minimize the epistemic uncertainty of the underlying model [11, 12, 13]. Model-based input signal design can, in principle, generate highly informative datasets by directly optimizing process-focused optimality criteria. However, the success of these approaches heavily relies on the employed model’s capability to accurately capture the process. Ensuring this capability in turn depends on prior knowledge – information that is often unavailable.
@@
novel contributionadaptive modelingcoarse modelAPRBS [14], multisine [15]Herkersdorf [5], Heinz [16] , Smits [17]Vater [18], Kiss [19], Liu [20]FIM-focused [7, 8, 9]uncertainty-focused [11, 12, 13]model-freemodel-based
@@ I-B Space-Filling Design in Model Input Space
The proposed method further distinguishes from existing approaches by introducing a novel optimality criterion embedded within a receding horizon control (RHC)-based optimization framework. In essence, the approach aims to generate a space-filling design within the relevant operating region of a surrogate model’s input space. This input space may correspond, for instance, to the joint input-state space, as in [18, 19, 20], or to the model regressor space, as considered in [5, 16, 17]. The fundamental premise is that the information an excitation signal provides about a process is inherently tied to the distribution it induces within the selected model space: if all relevant regions of a model’s input space are sufficiently represented in the estimation dataset and the chosen model accurately predicts the process, comprehensive information is acquired about the operating region of interest. This consideration is particularly true for Markovian processes, where future states can be fully predicted given the knowledge of the current state [21]. Consequently, the task of acquiring valuable information about the process behavior can be reframed as the challenge of sufficiently exciting the regions of interest within the model input space.
The generalization error, i.e., model errors in regions of the operating space not covered by the estimation data, is one of the main sources of inaccuracies in nonlinear system identification [1, 2]. A space-filling design within the operating region of interest ensures comprehensive data collection of the process behavior intended to be modeled, minimizing knowledge gaps. Moreover, evenly distributed data mitigates data bias toward specific local operating regions, reducing the risk of overfitting [22]. Consequently, space-filling excitation promotes the development of models that generalize well, contributing to robust and reliable predictions.
@@ II PROBLEM FORMULATION
$\hat{y}(k+1)=f\bigl{(}\underline{x}(k)\bigr{)}\hskip 3.0pt\quad\forall\,k\in\mathbb{N}\ ,$ (2)
@@
$\begin{split}{\underline{x}}^{T}(k+1)=&[u_{1}(k),\ldots,u_{1}(k-m+1),\ldots,\\ &u_{n_{u}}(k),\ldots,u_{n_{u}}(k-m+1),\\ &y_{1}(k),\ldots,y_{1}(k-m+1),\ldots,\\ &y_{n_{y}}(k),\ldots,y_{n_{y}}(k-m+1)]\quad\forall k\in\mathbb{N}\ ,\\ \end{split}$ (3)
@@
In a data-driven modeling framework, $f(\cdot)$ in (2) is inferred from estimation data, making its accuracy contingent on the quality of the available dataset. Hence, the general objective of input design is to generate a dataset
$\begin{split}{\underline{U}}^{T}=[\underline{u}(1),\ldots,\underline{u}(N)]\ ,\hskip 9.0ptN\in\mathbb{N}\end{split}$ (4)
of size $N$, where $\underline{u}^{T}(j)=[{u}_{1}(j),\ldots,{u}_{n_{u}}(j)]\ \forall j\in\{1,2,\ldots,N\}$, which should capture as much information as possible about the process behavior intended to be modeled. This ensures that $f(\cdot)$ in (2) can be estimated accurately, thereby minimizing the error specified in (1). In this contribution, the objective is pursued by optimizing the surrogate model’s input space distribution according to a space-filling optimality criterion. For an NARX structure, this distribution corresponds to the regressor space distribution, which can be expressed for an excitation dataset as
$\begin{split}\underline{{X}}^{T}=[{\underline{x}}(1),{\underline{x}}(2),\ldots,{\underline{x}}(N)]\ ,\hskip 9.0ptN\in\mathbb{N}\ \ .\end{split}$ (5)
The initial state of the process $\underline{x}(0)$, as well as the safe operating regions for the controllable inputs $\mathcal{U}$ and the regressor space $\mathcal{X}$, are assumed to be known, allowing the formulation of the following constraints to be satisfied:
$\begin{split}&\underline{{u}}(j)\in\mathcal{U},\ \underline{{x}}(j)\in\mathcal{X}\hskip 3.0pt\quad\forall\,j=\{1,2,\ldots,N\}\ .\end{split}$ (6)
@@ III RHC-BASED SPACE-FILLING DESIGN
In this section, the proposed input signal design methodology is introduced, and its key contributions are highlighted. First, we develop a receding horizon control (RHC)-based iterative optimization procedure, following a similar approach to [23, 24]. Second, we formulate a novel optimality criterion that is embedded in the RHC-strategy and optimizes a surrogate model’s input space distribution to achieve a space-filling design within the operating region of interest, thus ensuring comprehensive and balanced information acquisition about the process behavior intended to be modeled.
@@ III-A Novel Optimality Criterion
$\begin{split}\underline{{\psi}}(j)\in\mathcal{C}\hskip 3.0pt\quad&\forall j={1,2,\ldots,N_{\Psi}},\hskip 9.0ptN_{\Psi}\in\mathbb{N}\ ,\end{split}$ (7)
@@
$\begin{split}&J\bigl{(}\underline{\hat{X}}(k),\underline{{\Psi}}\bigr{)}=\sum_{j=1}^{N_{\Psi}}d_{\mathrm{NN}}\bigl{(}{\underline{\psi}}(j),\underline{\hat{X}}(k)\bigr{)}\\ &\mathrm{with\,}\,d_{\mathrm{NN}}\bigl{(}{\underline{\psi}}(j),\underline{\hat{X}}(k)\bigr{)}=\underset{1\leq i\leq k+L-1}{\min}||\underline{\hat{x}}(i)-\underline{\psi}(j)||_{2}\,.\end{split}$ (8)
An NARX structure is selected as the model input space; therefore, the model input space distribution corresponds to the regressor space distribution $\underline{\hat{X}}(k)$, approximated by the surrogate model at time step $k$ of the signal. This approximation is necessary since access to future outputs – required for executing the RHC-based optimization procedure, see Sec. III-B – as well as past outputs when running the proposed algorithm offline, is essential but unavailable. Therefore, the yet unknown true process outputs $y$ are substituted with the surrogate model’s predictions $\hat{y}$.
@@ III-B RHC-Based Optimization Procedure
The core principle of the RHC-based procedure lies in its iterative optimization. In each iteration, the optimal sequence $\tilde{\underline{U}}_{\mathrm{opt}}(k)$, starting at the current time step $k$ over a finite prediction horizon $L$, is determined as follows:
$\begin{split}&\underline{\tilde{U}}_{\mathrm{opt}}(k)=\underset{\underline{\tilde{U}}}{\mathrm{arg\,min}}\ J\bigl{(}\underline{\hat{X}}(k),\underline{{\Psi}}\bigr{)}\\ \mathrm{with}\ &\underline{\hat{X}}(k)=\mathcal{M}_{\theta}([\underline{U},\underline{\tilde{U}}]^{T},\underline{\hat{x}}(0))\\ \mathrm{s.\,t.\ }\ &\underline{\tilde{u}}(j)\in\mathcal{U}\hskip 3.0pt\quad\forall\,j=\{k,\ldots,k+L-1\}\\ &\underline{\hat{x}}(j)\in\mathcal{\mathcal{X}}\hskip 3.0pt\quad\forall\,j=\{k,\ldots,k+L-1\}\ .\end{split}$ (9)
Here, $\mathcal{M}_{\theta}$ represents the surrogate model, $\underline{U}$ denotes the previously optimized input data, and $\underline{\tilde{U}}$ refers to data points yet to be optimized. For a first-order dynamic system ($m=1$) with a single input ($n_{u}=1$) and a single output ($n_{y}=1$), the regressor space distribution constructed using the surrogate model at time step $k$ is defined by
$\underline{\hat{X}}(k)$ $=\begin{bmatrix}u(1)&\hat{y}(1)\\ \vdots&\vdots\\ u(k-1)&\hat{y}(k-1)\\ \tilde{u}(k)&\hat{y}(k)\\ \vdots&\vdots\\ \tilde{u}(k+L-1)&\hat{y}(k+L-1)\end{bmatrix}\ ,$ (10)
@@
Once $\underline{\tilde{U}}_{\mathrm{opt}}(k)$ is determined, its first data point $\tilde{\underline{u}}_{\mathrm{opt}}(k)$ is concatenated with the already optimized input data $\underline{U}$, and a forward time shift is performed, updating $k\rightarrow k+1$. This optimization procedure is iteratively repeated until the entire signal is generated, i.e., $k=N$. Pseudocode for the proposed procedure is presented in Algorithm 1.
@@
Initialization: Safe operating region of the controllable inputs $\mathcal{U}$, safe operating region of the regressor space $\mathcal{X}$, supporting points $\underline{\Psi}$, surrogate model $\mathcal{M}_{\theta}$, initial state $\underline{\hat{x}}(0)$.
@@
  $\hskip 0.0pt\underline{\tilde{U}}_{\mathrm{opt}}(k)=\underset{\underline{\tilde{U}}}{\mathrm{arg\,min}}\ J\bigl{(}\underline{\hat{X}}(k),\underline{{\Psi}}\bigr{)}$triangleright Get optimal xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxsequence.
$\begin{split}\mathrm{with}\ &\underline{\hat{X}}(k)=\mathcal{M}_{\theta}([\underline{{U}},\underline{\tilde{U}}]^{T},\underline{\hat{x}}(0))\\ \mathrm{s.\,t.\ }\ &\underline{\tilde{u}}(j)\in\mathcal{U}\hskip 3.0pt\quad\forall\,j=\{k,\ldots,k+L-1\}\\ &\underline{\hat{x}}(j)\in\mathcal{\mathcal{X}}\hskip 3.0pt\quad\forall\,j=\{k,\ldots,k+L-1\}\\ \end{split}$
  Apply $\underline{U}\leftarrow\underline{{U}}\cup\underline{\tilde{u}}_{\mathrm{opt}}(k)$. $\triangleright$ Append optimal xxxxxxxxxxxxxxxxxxxxxxxxxxxxxxxx data point.
@@
Figure 2 provides an illustrative example of the optimization process at a given time step $k$. The previously optimized data are indicated by  , while the optimal future sequence is highlighted in  . Together, these form the regressor space distribution $\underline{{\hat{X}}}(k)$. Additionally, the uniformly distributed supporting dataset $\underline{\Psi}$, generated using Sobol sequences [25], is depicted. To enhance visualization $N_{\Psi}<N$ is presented. In practice, however, it is advisable to choose the number of supporting points such that $N_{\Psi}\gg N$, allowing gaps to be efficiently targeted. By default, $N_{\Psi}=5N$.
[…]
Figure 2: Illustrative example of the optimization process at time step $k$: (a) input and corresponding output values, with previously optimized data in  , and the optimal future sequence in  . (b) distribution of the regressor space $\underline{\hat{X}}(k)$ as approximated by the surrogate model in $\circ$ / $\circ$, together with the supporting dataset $\underline{{\Psi}}$ in $\bullet$.
III-C Implementation Details and Execution Modes
Constraining the controllable inputs and the surrogate regressor space, as outlined in  (LABEL:eq:constraints), is particularly beneficial for real-world applications, as it prevents operation in undesirable regions.
In this contribution, the optimization focuses on the surrogate regressor space distribution of the generated excitation signal. However, optimizing the joint input-state space obtained by a surrogate, as explored in [18, 19, 20], is also a viable approach.
@@ IV EVALUATION
Unfortunately, the concept of space-filling design is not well-defined, particularly for dynamic processes. However, two key aspects are generally considered important: (i) ensuring that no region remains underrepresented, meaning that no gaps are left [26], and (ii) achieving a distribution that covers the considered area as uniformly as possible [27]. Hence, to evaluate the space-filling properties, two metrics are used. First, we examine the radius of the largest empty ball within the region of interest $\mathcal{C}$, defined as
@@
where $X$ denotes the data’s probability distribution, $E$ the uniform distribution, ${M}=\frac{1}{2}({X}+{E})$ the mixture distribution, and $D_{\mathrm{KL}}$ the Kullback-Leibler divergence [28].
Furthermore, both execution approaches presented–online execution with continuous updates of the model parameters based on current process observations, and offline execution with a fixed model, utilizing only an approximate process time constant as in [5, 16, 17]–are considered.
A Hammerstein process is employed as example process in the evaluation, defined by: